Make a quadratic equation whose roots have a sum of 4 and a product of 3 and another equation for the sum of 16 and a product of -80.
Answer by edjones(8007)

x+y=4
xy=3
y=4-x
x(4-x)=3
4x-x^2=3
x^2-4x+3=0
.
Ed